OpenStudy (anonymous):

What is the volume of the cube with side length seventeen inches? two hundred four cubic inches four hundred eight cubic inches one thousand seven hundred thirty-four cubic inches four thousand nine hundred thirteen cubic inches @bradely @dg2

OpenStudy (anonymous):

@hello_j

OpenStudy (bradely):

Volume of the cube:V= a^3 a =17 V= 17^3 =4913 Option 4 is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Can you help me with a few more?